# Loss Gradient Gaussian Width based Generalization and Optimization Guarantees

### v1

@@ B.1 GD with sample reuse
$\mathbb{P}_{z^{(n)}\sim{\cal D}^{n}}\left[\max_{t\in[T]}X_{t}>\epsilon\right]$ $=\mathbb{P}_{{z^{(n)}\sim{\cal D}^{n}}}\left[\max_{t\in[T]}\sup_{\theta_{t}\in\Theta}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}>\epsilon\right]~{}.$
From Lemma 3, for any $\delta>0$, with probability at least $1-\delta$ over the data $z^{(n)}$, we have
Choosing $u=\sqrt{(\log T+\log\frac{1}{\delta})/2}$ and $\epsilon=\frac{c_{1}w(\hat{\Xi}_{n})+u}{\sqrt{n}}$, we have
$\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T]}X_{t}>\epsilon\right]$ $\leq\sum_{t=1}^{T}\mathbb{P}_{z^{(n)}}\left[X_{t}>\epsilon\right]$
$\overset{(a)}{\leq}T\exp\left(-\left(\log T+\log\frac{1}{\delta}\right)\right)$
$=\delta,$ (58)
where $(a)$ follows by union bound and taking $\theta_{t}$ in $X_{t}$ as $\theta$ in (57). Then, for $\delta\leq 1/2$, with
$\epsilon=\frac{c_{1}w(\hat{\Xi}_{n})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{n}}$ (59)
we have
$\mathbb{P}_{{\cal D}_{T}(z^{(n)})}\left[\max_{t\in[T]}X_{t}>\epsilon\right]$ $\leq\frac{\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T]}X_{t}>\epsilon\right]}{\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T]}X_{t}\leq\epsilon\right]}\leq\frac{\delta}{1-\delta}\leq 2\delta~{}.$
In other words, with probability at least $(1-2\delta)$, for adaptively chosen $\theta_{t}$, we have
@@
$\max_{t\in[T]}\sup_{\theta_{t}\in\Theta}\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{{\cal D}}(\theta_{t})\right\|_{2}\leq\frac{c_{1}w(\hat{\Xi}_{n})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{n}}~{}.$
@@
For any $\delta>0$, with probability at least $1-\delta$ over the data $z^{(n)}$, with $\hat{\Xi}_{n},w(\hat{\Xi}_{n})$ respectively denoting the set of empirical gradients and its Gaussian width as in Definition 2, for any $u>0$, we have
@@
B.2 Mini-Batch SGD with Sample Resuse
Mini-batch Stochastic Gradient Descent. For mini-batch SGD, we consider a mini-batch size of $b\leq n$ and a total of $T$ update steps, yielding parameters $\theta_{t},t=1,\ldots,T$. For convenience, we assume $n/b$, the number of mini-batches in each epoch, and $K=T/n$, the total number of epochs, are integers. In the first epoch $\kappa=1$, we can split $z^{(n)}$ into $n/b$ groups of $b$ fresh samples each, and get $\theta_{1}$ using group $h=1$, $\theta_{2}$ using group $h=2$, and so on till we update $\theta_{n/b}$ using group $h=n/b$. For all subsequent epochs $\kappa=2,\ldots,K$, we can similarly do $n/b$ mini-batch updates based on the draw of the mini-batches or, as often used in practice, by permuting $z^{(n)}$ and picking minibatches sequentially [68, 76, 35]. Our results do not depend on how the mini-batches are chosen. For convenience, we will use $\theta_{(\kappa,h)}:=\theta_{(\kappa-1)\frac{n}{b}+h}$, i.e., parameter in epoch $\kappa$ and step $h$, so that $\kappa\in[K],h\in[n/b]$ indexes all the $T$ parameters. Consider the single gradient set for $\theta\in\Theta$:
$\hat{\Xi}(\Theta):=\left\{\hat{\xi}(\theta,z)\in\mathbb{R}^{p}\mid\exists\theta\in\Theta,z\in z^{(n)}~{}~{}~{}\text{s.t.}~{}~{}\hat{\xi}(\theta,z)=\nabla\ell(\theta;z),\right\}~{}.$ (65)
Denoting $\hat{\Xi}(\Theta)=\hat{\Xi}$ for simplification, we have the following result:
Theorem 7.
Let $\theta_{(\kappa,h)}\in\Theta$ be a sequence of parameters obtained with $T=K\frac{n}{b}$ steps of mini-batch SGD with batch-size $b$ over epochs $\kappa\in[K]$ and steps $h\in[\frac{n}{b}]$ in each epoch, by reusing a fixed set of samples $z^{(n)}\sim{\cal D}^{n}$ in each epoch $\kappa\in[K]$. Then, for any $\delta\leq\frac{1}{2}$, with probability at least $(1-2\delta)$ over $z^{(n)}\sim{\cal D}^{n}$, we have
$\max_{\begin{subarray}{c}\kappa\in[K]\\ h\in[\frac{n}{b}]\end{subarray}}\sup_{\theta_{(\kappa,h)}\in\Theta}\left\|\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)};z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\right\|_{2}\leq\frac{c_{1}w(\hat{\Xi})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}~{}.$
@@ B.2 Mini-Batch SGD with Sample Resuse
The proof is essentially a variant of the proof of Theorem 2 while make suitable changes due to the focus on mini-batches of $b$ samples. The goal is to make sure that for an adaptively chosen sequence of functions $f_{t},t=1,\ldots,T$, the deviation
$\max_{\kappa\in[K]}\max_{h\in[\frac{n}{b}]}\sup_{\theta_{(\kappa,h)}\in\Theta}\left\|\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\right\|_{2}$ (66)
is small. Note that (66) considers the largest deviation between the mini-batch and population expectation over all the $Kn/b=T$ vector-valued functions. Following the proof of Theorem 2, it suffices to first focus on uniformly bounding the deviation for each event in the non-adaptive setting, extend the argument to all $T$ events using union bound, and use the fact that the probability of the adaptive event can be upper bounded by the ratio of two non-adaptive events as in (56).
Let $z^{(b)}$ be the mini-batch drawn from $z^{(n)}$. We denote the set of all possible mini-batch average gradients as
$\hat{\Xi}_{n}^{b}:=\{\hat{\xi}_{b}\in\mathbb{R}^{p}\mid\exists\theta\in\Theta,z^{(b)}\subseteq z^{(n)}\text{s.t.}~{}~{}\hat{\xi}_{b}=\frac{1}{b}\sum_{z\in z^{(b)}}\nabla\ell(\theta;z)\}~{}.$ (67)
Note that when $b=n$, i.e., full batch, we have $\hat{\Xi}_{n}^{b}=\hat{\Xi}_{n}$.
At each iteration, for a fixed mini-batch $z_{t}^{(b)}$ drawn from fixed set $z^{(n)}$, consider the mini-batch average gradient
$\hat{\Xi}_{n}^{b,t}(\Theta):=\left\{\hat{\xi}_{b}\in\mathbb{R}^{p}\mid\exists\theta\in\Theta,~{}~{}~{}\text{s.t.}~{}~{}\hat{\xi}_{b}=\frac{1}{b}\sum_{z\in z_{t}^{(b)}}\nabla\ell(\theta;z),\right\}~{},$ (68)
We have $\hat{\Xi}_{n}^{b,t}(\Theta)\subseteq\hat{\Xi}_{n}^{b}(\Theta)$ so that $w(\hat{\Xi}_{n}^{b,t}(\Theta))\leq w(\hat{\Xi}_{n}^{b}(\Theta))$. We denote $\hat{\Xi}_{n}^{b}$ as $\hat{\Xi}_{n}^{b}(\Theta)$ as before to simplify the notation. Further, similar to before, we use $\hat{\Xi}_{n}^{(b)}$ to denote the set of $b$-tuples of gradients for $z^{(b)}\subseteq z^{(n)}$, over all $z^{(b)}$ and $\theta\in\Theta$.
Then, with $\hat{R}_{b}(\hat{\Xi}^{(b)}_{n})$ denoting the NERC for $\hat{\Xi}_{n}^{(b)}$, and $\bar{d}^{(b)}_{n}$ denoting the equivalent non-canonical distance over the $b$-tuples, by a direct extension of the proof argument for Theorem 1 using the new notation, we get
$\hat{R}_{b}(\hat{\Xi}^{(b)}_{n})\leq\frac{c_{2}\gamma_{2}(\hat{\Xi}^{(b)}_{n},\bar{d}^{(b)}_{n})}{\sqrt{b}}\leq\frac{c_{3}\gamma_{2}(\hat{\Xi}_{n}^{b},d)}{\sqrt{b}}\leq\frac{c_{4}w(\hat{\Xi}_{n}^{b})}{\sqrt{b}}~{}.$ (69)
Then, by a straightforward extension of Lemma 3, we have
$\mathbb{P}_{z^{(b)}}\left[\sup_{\theta_{(\kappa,h)}\in\Theta}\bigg{\|}\frac{1}{b}\sum_{z_{j}\in z_{t}^{(b)}}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\bigg{\|}_{2}>c_{1}\frac{w(\hat{\Xi}_{n}^{b})+u}{\sqrt{b}}\right]\leq c_{0}\exp(-2u^{2})~{}.$ (70)
Then, choosing $u=\sqrt{(\log T+\log\frac{1}{\delta})/2}$ and recalling that $\kappa\in[K],h\in[n/b]$, we have
$\mathbb{P}_{z^{(b)}}$ $\left[\max_{\kappa,h}\sup_{\theta_{(\kappa,h)}\in\Theta}\left\|\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\right\|_{2}>\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}\right]$
$=\mathbb{P}_{z^{(b)}}\left[\max_{\kappa,h}\sup_{\theta_{(\kappa,h)}\in\Theta}\bigg{\|}\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\bigg{\|}_{2}>\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}\right]$
$\leq\sum_{\kappa,h}\mathbb{P}_{z^{(b)}}\left[\sup_{\theta_{(\kappa,h)}\in\Theta}\bigg{\|}\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\bigg{\|}_{2}>\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}\right]$
$\leq\sum_{\kappa,h}\mathbb{P}_{z^{(b)}}\left[\sup_{\theta_{(\kappa,h)}\in\Theta}\bigg{\|}\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\bigg{\|}_{2}>\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}\right]$
$\leq T\exp\left(-\left(\log T+\log\frac{1}{\delta}\right)\right)$
$=\delta~{}.$
Then, for $\delta\leq 1/2$, following notation in (56) for the GD case to discuss the adaptive event, with
$\epsilon_{0}=\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}$ (71)
we have
@@
$\mathbb{P}_{{\cal D}_{T}{(z^{(b)})}}$ $\left[\max_{\kappa,h}\sup_{\theta_{(\kappa,h)}\in\Theta}\left\|\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\right\|_{2}>\epsilon_{0}\right]$
@@
$\leq\frac{\mathbb{P}_{z^{(b)}}\left[\max_{\kappa,h}\sup_{\theta_{(\kappa,h)}\in\Theta}\left\|\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\right\|_{2}>\varepsilon_{0}\right]}{\mathbb{P}_{z^{(b)}}\left[\max_{\kappa,h}\sup_{\theta_{(\kappa,h)}\in\Theta}\left\|\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\right\|_{2}\leq\varepsilon_{0}\right]}$
$\leq\frac{\delta}{1-\delta}$
$\leq 2\delta~{}.$
In other words, with probability at least $(1-2\delta)$, for adaptively chosen $\theta_{(\kappa,h)}$, we have
@@
$\max_{\kappa,h}\sup_{\theta_{(\kappa,h)}\in\Theta}\left\|\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell(\theta_{(\kappa,h)},z_{j})-\nabla{\cal L}_{{\cal D}}(\theta_{(\kappa,h)})\right\|_{2}\leq\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}~{}.$ (72)
To generalize the result, we define the Gaussian width on the set of single sample gradients. Consider the single gradient set for $\theta\in\Theta$:
@@
$\hat{\Xi}(\Theta):=\left\{\hat{\xi}(\theta,z)\in\mathbb{R}^{p}\mid\exists\theta\in\Theta,z\in z^{(n)}~{}~{}~{}\text{s.t.}~{}~{}\hat{\xi}(\theta,z)=\nabla\ell(\theta;z),\right\}~{}.$ (73)
Denote $\hat{\Xi}(\Theta)$ as $\hat{\Xi}$, we have the Gaussian width on the single gradient set as
$w(\hat{\Xi})$ $=\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta\in\Theta,z\in z^{(n)}}~{}\langle\hat{\xi}(\theta,z),\mathbf{g}\rangle\right]$ (74)
We have the Gaussian width on the mini-batch gradient set as
@@
$w(\hat{\Xi}_{n}^{b})$ $=\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta\in\Theta,z^{(b)}\subseteq z^{(n)}}~{}\left\langle\frac{1}{b}\sum_{z\in z^{(b)}}\hat{\xi}(\theta,z),~{}\mathbf{g}\right\rangle\right]$
@@
$\leq\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta\in\Theta,z^{(b)}\subseteq z^{(n)}}~{}\left\langle\frac{1}{b}\sum_{z\in z^{(b)}}\hat{\xi}(\theta,z),~{}\mathbf{g}\right\rangle\right]$
@@
$=\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta\in\Theta,z^{(b)}\subseteq z^{(n)}}~{}\frac{1}{b}\sum_{z\in z^{(b)}}\left\langle\hat{\xi}(\theta,z),~{}\mathbf{g}\right\rangle\right]$
$\leq\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta\in\Theta,z\in z^{(b)},z^{(b)}\subseteq z^{(n)}}~{}\left\langle\hat{\xi}(\theta,z),~{}\mathbf{g}\right\rangle\right]$
$=\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta\in\Theta,z\in z^{(n)}}~{}\langle\hat{\xi}(\theta,z),\mathbf{g}\rangle\right]$
$=w(\hat{\Xi})$ (75)
Plugging in $w(\hat{\Xi})$ to Equation 72 completes the proof. ∎
B.3 Population Convergence in Optimization

### v2

@@ B.1 GD with sample reuse
$\mathbb{P}_{z^{(n)}}\left[\sup_{\theta_{0}\in\Theta_{0}}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta,z_{i})-\nabla{\cal L}_{{\cal D}}(\theta)\bigg{\|}_{2}>\frac{c_{1}w(\hat{\Xi}^{0}_{n})+u}{\sqrt{n}}\right]\leq\exp(-2u^{2})~{}.$ (61)
Further, from Lemma 5, for any $u>0$, we have
@@
$\mathbb{P}_{z^{(n)}}\left[\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta,z_{i})-\nabla{\cal L}_{{\cal D}}(\theta)\bigg{\|}_{2}>\frac{c_{1}\left(\log p+u\right)}{\sqrt{n}}\right]\leq\exp(-u^{2})~{}.$ (62)
Choosing $\epsilon=\frac{c_{1}\max(w(\hat{\Xi}_{n}),\log p)+u}{\sqrt{n}}$ and $u=\log(T+1)+\log\frac{1}{\delta}$, we have
$\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T\}}X_{t}>\epsilon\right]$ $\overset{(a)}{\leq}\sum_{t=0}^{T}\mathbb{P}_{z^{(n)}}\left[X_{t}>\epsilon\right]=\mathbb{P}_{z^{(n)}}\left[X_{0}>\epsilon\right]+\sum_{t=1}^{T}\mathbb{P}_{z^{(n)}}\left[X_{t}>\epsilon\right]$
$\overset{(b)}{\leq}\exp(-2u^{2})+T\exp(-u^{2})$
$\overset{(c)}{\leq}(T+1)\exp(-u^{2})$
$\overset{(d)}{\leq}(T+1)\exp\left(-\left(\log(T+1)+\log\frac{1}{\delta}\right)\right)=\delta,$
where $(a)$ follows by union bound, (b) follows from 61 and 62, (c) follows since $u>1$ for $T\geq 2$, and (d) follows from the form of $u$. Then, for $\delta\leq 1/2$ we have
$\mathbb{P}_{{\cal D}_{T}(z^{(n)})}\left[\max_{t\in\left\{0:T\right\}}X_{t}>\epsilon\right]$ $\leq\frac{\mathbb{P}_{z^{(n)}}\left[\max_{t\in\left\{0:T\right\}}X_{t}>\epsilon\right]}{\mathbb{P}_{z^{(n)}}\left[\max_{t\in\left\{0:T\right\}}X_{t}\leq\epsilon\right]}\leq\frac{\delta}{1-\delta}\leq 2\delta~{}.$
@@
With $\hat{\Xi}_{n},w(\hat{\Xi}_{n})$ respectively denoting the set of empirical gradients and its Gaussian width as in Definition 2, for any $u>0$, we have
@@
Lemma 4 (Matrix Bernstein Inequality).
[5] Let $S_{1}$,…,$S_{n}$ be independent, centered random matrices with common dimension $d_{1}\times d_{2}$, and assume that each one is uniformly bounded
$\mathbb{E}S_{k}=0,\left\|S_{k}\right\|\leq L,\forall k\in[n].$
Introduce the sum
$Z=\sum_{k=1}^{n}S_{k},$
and let $v(Z)$ denote the matrix variance statistic of the sum
$v(Z)$ $=\max\left\{\left\|\mathbb{E}\left(ZZ^{*}\right)\right\|,\left\|\mathbb{E}\left(Z^{*}Z\right)\right\|\right\}$
$=\max\left\{\left\|\sum_{k=1}^{n}\mathbb{E}\left(S_{k}S_{k}^{*}\right)\right\|,\left\|\sum_{k=1}^{n}\mathbb{E}\left(S_{k}^{*}S_{k}\right)\right\|\right\}.$
Then
$\mathbb{P}\left\{\left\|Z\right\|\geq t\right\}\leq\left(d_{1}+d_{2}\right)\cdot\exp\left(\frac{-t^{2}/2}{v(Z)+Lt/3}\right),\forall t\geq 0.$
Furthermore,
$\mathbb{E}\left\|Z\right\|\leq\sqrt{2v(Z)\log(d_{1}+d_{2})}+\frac{1}{3}L\log(d_{1}+d_{2}).$
Lemma 5.
Under Assumption 2, for any $u>0$, we have
$\mathbb{P}_{z^{(n)}}\left[\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta,z_{i})-\nabla{\cal L}_{{\cal D}}(\theta)\bigg{\|}_{2}>\frac{c_{1}\left(\log p+u\right)}{\sqrt{n}}\right]\leq\exp(-u^{2})~{}.$ (68)
@@
Following Assumption 2, denote the $L_{2}$ norm bound of the gradient is $G$. There are two different cases:
  • If $\log p+u\geq\sqrt{n}$, then applying the bounded difference inequality [16], for any $\epsilon>0$,
@@
    $\mathbb{P}_{z^{(n)}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell\left(\theta,z_{i}\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|_{2}>\mathbb{E}_{z^{(n)}}\left\|\nabla\hat{\mathcal{L}}_{n}\left(\theta\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|+\epsilon\right]\leq\exp\left(-\frac{n\epsilon^{2}}{2G^{2}}\right)$
    Since
@@
    $\mathbb{E}_{z^{(n)}}\left\|\nabla\hat{\mathcal{L}}_{n}\left(\theta\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|\leq 2G=\frac{2\sqrt{n}G}{\sqrt{n}}\leq\frac{2G\left(\log p+u\right)}{\sqrt{n}}$
    then taking $\epsilon=\frac{\sqrt{2}Gu}{\sqrt{n}}$, we have
@@
    $\mathbb{P}_{z^{(n)}}$ $\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell\left(\theta,z_{i}\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|_{2}>\frac{2G\left(\log p+u\right)+\sqrt{2}Gu}{\sqrt{n}}\right]$
@@
    $\leq\mathbb{P}_{z^{(n)}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell\left(\theta,z_{i}\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|_{2}>\mathbb{E}_{z^{(n)}}\left\|\nabla\hat{\mathcal{L}}_{n}\left(\theta\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|+\frac{\sqrt{2}Gu}{\sqrt{n}}\right]$
    $\leq\exp(-u^{2})$
    Therefore,
@@
    $\mathbb{P}_{z^{(n)}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell\left(\theta,z_{i}\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|_{2}>\frac{2G\left(\log p+u\right)+\sqrt{2}Gu}{\sqrt{n}}\right]\leq\exp(-u^{2})$
  • If $\log p+u<\sqrt{n}$, then with $S_{i}=\nabla\ell\left(\theta_{t},z_{i}\right)-\mathbb{E}\nabla\ell\left(\theta_{t},z_{i}\right)$ and $Z=\sum_{i=1}^{n}S_{i}$, we have $\mathbb{E}S_{i}=0$, and
    $\left\|S_{i}\right\|=\left\|\nabla\ell\left(\theta_{t},z_{i}\right)-\mathbb{E}\nabla\ell\left(\theta_{t},z_{i}\right)\right\|\leq\left\|\nabla\ell\left(\theta_{t},z_{i}\right)\right\|+\left\|\mathbb{E}\nabla\ell\left(\theta_{t},z_{i}\right)\right\|\leq 2G.$
    so we can take $L=2G$ as an upper bound of $\left\|S_{i}\right\|$. Besides,
    $v(Z)=\max\left\{\left\|\sum_{i=1}^{n}\mathbb{E}\left(S_{i}S_{i}^{*}\right)\right\|,\left\|\sum_{i=1}^{n}\mathbb{E}\left(S_{i}^{*}S_{i}\right)\right\|\right\}\leq n\max_{i\in[n]}\left\|S_{i}\right\|^{2}\leq 4nG^{2}$
    Then directly apply Lemma 4 with $d_{1}=1$, $d_{2}=p$, we can get that
    $\mathbb{P}\left\{\left\|Z\right\|\geq t\right\}\leq\left(p+1\right)\cdot\exp\left(-\frac{t^{2}/2}{v(Z)+Lt/3}\right)$
    Substituting $L=2G$ and $v(Z)\leq 4nG^{2}$, we have
    $\mathbb{P}\left\{\left\|Z\right\|\geq t\right\}\leq\left(p+1\right)\cdot\exp\left(-\frac{t^{2}/2}{4nG^{2}+2Gt/3}\right)=\left(p+1\right)\cdot\exp\left(\frac{-3t^{2}}{24nG^{2}+4Gt}\right)$
    Taking $t=4\sqrt{n}G\left(\log p+u\right)$, we have
    $\mathbb{P}\left\{\left\|Z\right\|\geq 4\sqrt{n}G\left(\log p+u\right)\right\}$ $\leq\left(p+1\right)\cdot\exp\left(-\frac{48nG^{2}\left(\log p+u\right)^{2}}{24nG^{2}+16\sqrt{n}G^{2}\left(\log p+u\right)}\right)$
    $<\left(p+1\right)\cdot\exp\left(-\frac{48nG^{2}\left(\log p+u\right)^{2}}{24nG^{2}+16\sqrt{n}G^{2}\cdot\sqrt{n}}\right)$
    $<\left(p+1\right)\cdot\exp\left(-\left(\log p+u\right)^{2}\right)<\left(p+1\right)\cdot\exp\left(-\left(\left(\log p\right)^{2}+u^{2}\right)\right)$
    $\leq\exp\left(-u^{2}\right)$
    So we have
@@
    $\mathbb{P}_{z^{(n)}}$ $\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell\left(\theta,z_{i}\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|_{2}>\frac{4G\left(\log p+u\right)}{\sqrt{n}}\right]$
    =$=\mathbb{P}\left\{\frac{1}{n}\left\|Z\right\|\geq\frac{4G\left(\log p+u\right)}{\sqrt{n}}\right\}=\mathbb{P}\left\{\left\|Z\right\|\geq 4\sqrt{n}G\left(\log p+u\right)\right\}$
    $\leq\exp(-u^{2})$
    Therefore,
@@
    $\mathbb{P}_{z^{(n)}}$ $\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell\left(\theta,z_{i}\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|_{2}>\frac{4G\left(\log p+u\right)}{\sqrt{n}}\right]\leq\exp(-u^{2})$
Combining these two cases, there exist $c_{3}$ such that
$\mathbb{P}_{z^{(n)}}$ $\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell\left(\theta,z_{i}\right)-\nabla\mathcal{L}_{\mathcal{D}}\left(\theta\right)\right\|_{2}>\frac{c_{3}G\left(\log p+u\right)}{\sqrt{n}}\right]\leq\exp(-u^{2})$
Since $G=O(1)$ according to Assumption 2, by taking $c_{1}=Gc_{3}$, we complete the proof.
B.2 Population Convergence in Optimization